# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

### v1

@@ Abstract
Li-CO2 batteries show promise as energy storage solutions, offering high theoretical energy density and CO2 fixation. Their performance relies on the formation and decomposition of Li2CO3/C during discharge and charge cycles, respectively. We used a multiscale modeling framework that integrates Density Functional Theory (DFT), Ab-Initio Molecular Dynamics (AIMD), classical Molecular Dynamics (MD), and Finite Element Analysis (FEA) to investigate atomic and cell-level properties. The Li-CO2 battery consists of a lithium metal anode, an ionic liquid electrolyte, and a carbon cloth porous cathode with a Sb0.67Bi1.33Te3 catalyst. DFT and AIMD determined the electrical conductivities of Sb0.67Bi1.33Te3 and Li2CO3 using the Kubo-Greenwood formalism and studied the CO2 reduction mechanism on the cathode catalyst. MD simulations calculated the CO2 diffusion coefficient, Li+ transference number, ionic conductivity, and Li+ solvation structure. The FEA model, incorporating results from atomistic simulations, reproduced experimental voltage-capacity profiles at 1 mA/cm2 and revealed spatio-temporal variations in Li2CO3/C deposition, porosity, and CO2 concentration dependence on discharge rates in the cathode. Accordingly, Li2CO3 can form toroidal and thin film deposits, leading to dispersed and local porosity changes at 0.1 mA/cm2 and 1 mA/cm2, respectively. The capacity decreases exponentially from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, due to pore clogging that limits CO2 transport to the cathode interior. Therefore, the performance of Li-CO2 batteries can be significantly improved by enhancing CO2 transport, regulating Li2CO3 deposition, and optimizing cathode architecture.
@@ 1 Introduction
Lithium-carbon dioxide (Li-CO2) batteries, with their high theoretical energy density and CO2 fixation capability, have emerged as a potential energy storage solution [1, 2]. These batteries provide a theoretical specific energy density of 1876 Wh kg-1, which is substantially higher than that of current lithium-ion battery systems [3, 4]. This high energy density makes Li-CO2 batteries particularly attractive for applications in aviation, aerospace, and electric vehicles, where lighter and more compact energy storage solutions are crucial [5]. The potential for electric vehicles with much longer ranges or considerably lighter batteries could revolutionize the automotive market [6, 5]. However, Li-CO2 batteries are still in early development stages and face challenges including reaction reversibility issues, low energy efficiency, and the need for improved cathode designs [7]. Ongoing research and development efforts focus on overcoming these limitations, with recent advancements showing promising results [4, 8, 9, 10, 11, 12, 13, 14]. The Li-CO2 batteries operate on a unique electrochemical principle involving the reversible reaction between lithium and carbon dioxide [1, 2] contributing to CO2 capture and utilization efforts [11, 15]. During the discharge process, lithium oxidizes at the anode, releasing lithium ions and electrons . Besides, at the cathode in the presence of solid catalysts [16, 4], CO2 undergoes reduction (), combining with lithium ions and electrons to form primarily lithium carbonate (Li2CO3) and carbon (C) . This discharge reaction typically occurs at a voltage of approximately 2.8 V [17]. The charging process reverses these reactions, decomposing the discharge products to regenerate lithium and CO2. However, due to the stability of Li2CO3, this process requires higher voltages, often exceeding 3.8 V [18]. This voltage requirement highlights one of several key challenges facing Li-CO2 battery technology. The formation of solid Li2CO3 and carbon can lead to electrode clogging and increased internal resistance [9, 19]. The high charging voltages necessary for Li2CO3 decomposition reduce overall energy efficiency [20, 21]. Effective cathode catalysts are critical for enhancing reaction kinetics and minimizing overpotentials in Li-CO2 batteries [8, 22]. Unwanted side reactions, particularly electrolyte decomposition, can occur at high voltages [7]. Furthermore, the inherent inertness of CO2 makes its electrochemical reduction challenging, affecting discharge voltage and efficiency [23, 2, 15].
@@
To bridge the gap between atomic and cell-level simulations, we first employ DFT/AIMD calculations to determine the electrical conductivity of Li2CO3 and Sb0.67Bi1.33Te3 using the Kubo-Greenwood formula [48, 49, 50], as well as to study the surface reaction of Li2CO3 as cathode catalyst material. Secondly, we conducted large-scale MD simulations to investigate the transport properties of the liquid electrolyte. Specifically, we calculate the diffusion coefficients of Li+ and CO2, and the Li+ transference number by examining the mean square displacement (MSD) of the particles over time using the Einstein equation [51], and the ionic conductivity using the Green-Kubo formula [52]. Finally, the atomic-scale insights gained from DFT/AIMD and MD simulations are then integrated into our FEA model, allowing for a comprehensive analysis of cell-level performance. Our FEA simulations investigate various aspects of battery cell behavior, including specific capacity, discharge behavior, and voltage-capacity profiles considering factors such as CO2 transport properties, cathode porosity, and deposition of Li2CO3 and carbon. The studied Li-CO2 battery cell consists of a lithium anode, an ionic liquid electrolyte, and a porous cathode. The electrolyte consists of a mixture of 1-ethyl-3-methylimidazolium tetrafluoroborate (EMIM-BF4) and dimethyl sulfoxide (DMSO) in a 2:3 ratio, supplemented with zinc iodide (ZnI2) and 1 M lithium bis(trifluoromethanesulfonyl)imide (LiTFSI) salt. This specific electrolyte composition has been experimentally demonstrated to yield optimal battery cell performance [4]. The cathode uses carbon cloth functionalized with a stable transition metal trichalcogenide alloy catalyst (Sb0.67Bi1.33Te3), improving CO2 reduction and oxidation kinetics [4]. This multiscale approach enables us to make predictions about porosity, deposited species quantities, and CO2 concentration variations-parameters that are difficult to assess through experimental methods alone. Furthermore, it allows us to evaluate these properties as a function of time, providing dynamic insights into the battery’s behavior. This computational methodology can be used to test and screen various battery materials, potentially saving significant time and resources by reducing the need for extensive experimental tests.
@@ 2.2 Ab initio molecular dynamics (AIMD) and Kubo-Greenwood method for electrical conductivity calculation
where is broadening of the -function. This technical parameter plays an essential role in controlling the algorithm in the Kubo-Greenwood formula, particularly at low frequencies.
@@ 2.4.1 Electrode Reactions
The main reaction at the cathode is described by [11]:
where and represent the equilibrium potentials for discharge and charge reactions, respectively [17, 18].
This reaction produces Li2CO3/C as the main products during discharge, which plays a crucial role in the battery’s performance. Li2CO3 and carbon are insoluble in the organic electrolyte and can cover the active surface area inside the porous cathode, block pathways for reactive species (Li+ and CO2), prevent further reactions inside the cathode and thus contribute to the end of cell discharge.
@@ 2.4.3 Equations of battery electrochemical properties
denotes the starting porosity of the cathode. It is important to mention that the effect of deposits (Li2CO3 and carbon) with low conductivity on was not taken into account, and it was assumed that would not change even as the battery reactions occurred.
@@
the Butler-Volmer equation is applied in the model using two rate coefficients because the reactions depends on both concentration of and and concentration of /C:
where is local transfer current density between electrode and
electrolyte interface, is the molar concentration of species at the cathode surface, and are the anodic and cathodic rate constant, respectively, is a symmetry factor equal to 0.5, is surface or activated overpotential for reaction at cathode, and are the voltage drop and the electrical resistivity across /C film formation, respectively, is the volume fraction of solid /C, and is the theoretical open-circuit potential for reaction.
@@ 3.1 Electrical conductivity from DFT, AIMD and Kubo-Greenwood formula
Calculated dynamic electrical conductivities for both Sb0.67Bi1.33Te3 and Li2CO3 are shown in Figure 2, for decremented values of from 0.9 to 0.54 with a step of 0.04. The electrical conductivities calculated for different values of show the conductive behavior of Sb0.67Bi1.33Te3 (a conductivity range of the order of S/m) and insulative behavior of Li2CO3 (a conductivity range of the order of S/m). Quantitatively, the electrical conductivities converge with decreasing , i.e., the dependence of conductivity on is smaller. The static electrical conductivities at zero frequency are calculated by interpolating small frequency data shown in Figure 2 (a2) and (b2). Using values where conductivity is converged, below 0.66 for Sb0.67Bi1.33Te3 and below 0.86 for Li2CO3, our computational analysis reveals a static electrical conductivity of S/m for Sb0.67Bi1.33Te3, indicating its strong electrical conducting properties. This finding aligns with the well-established characteristics of bismuth telluride (Bi2Te3) and its alloys, which are extensively employed in thermoelectric devices due to their good electrical conductivity [83, 84]. The high conductivity can be ascribed to the semiconductor nature of Bi2Te3 and the antimony (Sb) doping effect, which augments the carrier concentration [85]. Experimental investigations on Bi2Te3 have reported electrical conductivity values as high as S/m [84], further corroborating the computational results. Our calculations reveal that lithium carbonate (Li2CO3) exhibits a static electrical conductivity range of S/m, indicating its poor electrical conductivity and insulating behavior [86]. This calculated value aligns closely with experimentally reported values, S/m [87] and S/m [88]. The insulating nature of Li2CO3 can be attributed to its ionic nature and the strong ionic bonds that hinder electron mobility [89].
@@
Calculated dynamic electrical conductivities for both Sb0.67Bi1.33Te3 and Li2CO3 are shown in Figure 2 (a1) and (b1) for a large frequency range and in Figures 2 (a2) and (b2) for a small frequency range, for decremented values of from 0.9 to 0.54 with a step of 0.04. Figures 2 (a2) and (b2), the electrical conductivities converge, for small frequencies, with decreasing , i.e., the dependence of conductivity on becomes smaller. The static electrical conductivities at zero frequency are calculated by interpolating small frequency data shown in Figures 2 (a2) and (b2), using values where conductivity has converged, below 0.66 for Sb0.67Bi1.33Te3 and below 0.86 for Li2CO3. The static electrical conductivities calculated for different values of show the conductive behavior of Sb0.67Bi1.33Te3 (a conductivity range on the order of S/m) and insulative behavior of Li2CO3 (a conductivity range on the order of - S/m). Our computational analysis reveals a static electrical conductivity of S/m for Sb0.67Bi1.33Te3, indicating its strong electrical conducting properties. This finding aligns with the well-established characteristics of bismuth telluride (Bi2Te3) and its alloys, which are extensively employed in thermoelectric devices due to their good electrical conductivity [83, 84]. The high conductivity can be attributed to the semiconductor nature of Bi2Te3 and the antimony (Sb) doping effect, which augments the carrier concentration [85]. Experimental investigations on Bi2Te3 have reported electrical conductivity values as high as S/m [84], further corroborating the computational results. Our calculations reveal that lithium carbonate (Li2CO3) exhibits a static electrical conductivity range of S/m, indicating its poor electrical conductivity and insulating behavior [86]. This calculated value aligns closely with experimentally reported values of S/m [87] and S/m [88]. The insulating nature of Li2CO3 can be attributed to its ionic nature and the strong ionic bonds that hinder electron mobility [89]. In Li2CO3, Li+ ions act as the primary charge carriers rather than electrons, a characteristic typical of ionic compounds. The low conductivity of Li2CO3 presents challenges to battery performance, including high internal resistance and passivation of the electrode surface [90, 91]. To address these challenges, potential solutions include enhancing Li2CO3 conductivity, promoting its decomposition through conductive additives or catalysts, designing porous electrodes for increased surface area, and exploring alternative Li2CO3 structures with improved conductivity or decomposition properties.
@@
The DFT calculations of the reaction path for Li2CO3/C formation on the Sb0.67Bi1.33Te3 as Li-CO2 battery cathode catalyst have revealed intriguing insights into the adsorption behavior of reaction products and the step-wise mechanism of the discharge process. To comprehensively understand the formation of Li2CO3/C and C, we calculated the Gibbs free energy (G) changes for various elementary reaction steps on the Sb0.67Bi1.33Te3 catalyst surface, revealing a sequence of intermediate stages that elucidate the complex reaction pathway.
The Gibbs free energy of the reaction was calculated using the following equation:
@@
To level up the electrochemical step, a potential was applied to the formation of LiCO2. As can be seen in Figure 2 (c), this potential occurs at 2.3 V versus Li/Li+ electrode.
The reaction begins with the adsorption of CO2∗ on the Bi site, with a Bi-O distance of 3.5 Å. This initial step is marginally endergonic, with a G of 0.002 eV, indicating a slight energy barrier to CO2 adsorption. Following this step, the coupled addition of Li+ and an electron to the adsorbed CO2∗ forms the LiCO2∗ intermediate. This step is highly exergonic, with a G of -2.3 eV, suggesting a thermodynamically favorable process that likely drives the overall reaction forward.
The reaction pathway continues with the interaction between two LiCO2∗ intermediates, producing Li2CO2∗ and CO2∗. This step has a G of -1.2 eV, indicating another energetically favorable process. Subsequently, a pivotal step occurs: the conversion of Li2CO2∗ and CO2∗ into Li2CO3 and CO∗, with a G of 1.7 eV. This step is crucial as it marks the formation of one of the main discharge products, Li2CO3.
The final step in the pathway involves the generation of carbon, facilitated through the reaction of CO∗ and CO2∗. Interestingly, this step is energetically uphill, with a G of 2.8 eV, and also produces CO3∗ on the surface. This energy barrier for carbon formation could explain the observed preferential adsorption of carbon on the catalyst surface, as the system may need to overcome this energy barrier to complete the carbon formation process.
@@
These results highlight the essential role of coupled electron-cation transfer processes in the growth of discharge products. The strong exergonic nature of steps 2-4 suggests that these processes occur readily, while the endergonic nature of the final carbon formation step aligns with the observed preferential adsorption of carbon on the Sb0.67Bi1.33Te3 catalyst surface.
The lack of Li2CO3 adsorption on the catalyst surface, as indicated by the simulation, can be explained by the reaction pathway. Li2CO3 is formed in step 4, but this step also releases CO∗ and a free surface site (∗). This suggests that Li2CO3 formation occurs near, but not directly on, the catalyst surface, reducing the cathode porosity and possibly clogging the cathode pores.
@@
This section presents the results of MD simulations that elucidate the atomic-level dynamic and solvation properties of the Li-CO2 battery electrolyte depicted in Figure 1 (C). These simulations provide microscopic insights into the electrolyte’s behavior, bridging the gap between atomic and macroscopic transport properties in our multiscale approach.

### v5

@@ Abstract
Li-CO2 batteries show promise as energy storage solutions, offering high theoretical energy density and CO2 fixation. Their operation is based on the formation and decomposition of Li2CO3/C during discharge and charge cycles, respectively. We used a multiscale modeling framework that integrates Density Functional Theory (DFT), Ab-Initio Molecular Dynamics (AIMD), classical Molecular Dynamics (MD), and Finite Element Analysis (FEA) to investigate atomic and cell-level properties. The considered Li-CO2 battery consists of a lithium metal anode, an ionic liquid electrolyte, and a carbon cloth porous cathode with Sb0.67Bi1.33Te3 as a catalyst. DFT and AIMD determined the electrical conductivities of Sb0.67Bi1.33Te3 and Li2CO3 using the Kubo-Greenwood formalism and studied the CO2 reduction mechanism on the cathode catalyst. MD simulations calculated the CO2 diffusion coefficient, Li+ transference number, ionic conductivity, and Li+ solvation structure. The FEA model, incorporating results from atomistic simulations, reproduced experimental voltage-capacity profiles at 1 mA/cm2 and revealed spatio-temporal variations in Li2CO3/C deposition, porosity, and CO2 concentration dependence on discharge rates in the cathode. Accordingly, Li2CO3 can form large and thin film deposits, leading to dispersed and local porosity changes at 0.1 mA/cm2 and 1 mA/cm2, respectively. The capacity decreases exponentially from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, due to pore clogging from excessive discharge product deposition that limits CO2 transport to the cathode interior. Therefore, the performance of Li-CO2 batteries can be improved by enhancing CO2 transport, regulating Li2CO3 deposition, and optimizing cathode architecture.
@@ 1 Introduction
Lithium-carbon dioxide (Li-CO2) batteries, with their high theoretical energy density and CO2 fixation capability, have emerged as a potential energy storage solution [1, 2]. These batteries provide a theoretical specific energy density of 1876 Wh kg-1, which is substantially higher than that of current lithium-ion battery systems [3, 4]. This high energy density makes Li-CO2 batteries particularly attractive for applications in aviation, aerospace, and electric vehicles, where lighter and more compact energy storage solutions are crucial [5]. The potential for electric vehicles with much longer ranges or considerably lighter batteries could revolutionize the automotive market [6, 5]. However, Li-CO2 batteries are still in early development stages and face challenges including reaction reversibility issues, low energy efficiency, and the need for improved cathode designs [7]. Ongoing research and development efforts focus on overcoming these limitations, with recent advancements showing promising results [4, 8, 9, 10, 11, 12, 13, 14]. The Li-CO2 batteries operate on a unique electrochemical principle involving the reversible reaction between lithium and carbon dioxide [1, 2] contributing to CO2 capture and utilization efforts [11, 15]. During the discharge process, lithium oxidizes at the anode, releasing lithium ions and electrons . Besides, at the cathode in the presence of solid catalysts [16, 4], CO2 undergoes reduction (), combining with lithium ions and electrons to form primarily lithium carbonate (Li2CO3) and carbon (C) . This discharge reaction typically occurs at a voltage of approximately 2.8 V [17]. The charging process reverses these reactions, decomposing the discharge products to regenerate lithium and CO2. However, due to the stability of Li2CO3, this process requires higher voltages, often exceeding 3.8 V [18]. This voltage requirement highlights one of several key challenges facing Li-CO2 battery technology. The formation of solid Li2CO3 and C can lead to electrode clogging and increased internal resistance [9, 19]. The high charging voltages necessary for Li2CO3 decomposition reduce overall energy efficiency [20, 21]. Effective cathode catalysts are critical for enhancing reaction kinetics and minimizing overpotentials in Li-CO2 batteries [8, 22]. Unwanted side reactions, particularly electrolyte decomposition, can occur at high voltages [7]. Furthermore, the inherent inertness of CO2 makes its electrochemical reduction challenging, affecting discharge voltage and efficiency [23, 2, 15].
@@
To bridge the gap between atomic and cell-level simulations, we first employ DFT/AIMD calculations to determine the electrical conductivity of Li2CO3 and Sb0.67Bi1.33Te3 using the Kubo-Greenwood formula [48, 49, 50], as well as to study the surface reaction of Sb0.67Bi1.33Te3 as cathode catalyst material. Secondly, we conducted large-scale MD simulations to investigate the transport properties of the liquid electrolyte. Specifically, we calculate the diffusion coefficients of Li+ and CO2, and the Li+ transference number by examining the mean square displacement (MSD) of the particles over time using the Einstein equation [51], and the ionic conductivity using the Green-Kubo formula [52]. Finally, the atomic-scale insights gained from DFT/AIMD and MD simulations are then integrated into our FEA model, allowing for a comprehensive analysis of cell-level performance. Our FEA simulations investigate various aspects of battery cell behavior, including specific capacity, discharge behavior, and voltage-capacity profiles considering factors such as CO2 transport properties, cathode porosity, and deposition of Li2CO3 and C. The studied Li-CO2 battery cell consists of a lithium anode, an ionic liquid electrolyte, and a porous cathode. The electrolyte consists of a mixture of 1-ethyl-3-methylimidazolium tetrafluoroborate (EMIM-BF4) and dimethyl sulfoxide (DMSO) in a 2:3 ratio, supplemented with zinc iodide (ZnI2) and 1 M lithium bis(trifluoromethanesulfonyl)imide (LiTFSI) salt. This specific electrolyte composition has been experimentally demonstrated to yield optimal battery cell performance [4]. The cathode uses carbon cloth coated with a stable transition metal trichalcogenide alloy catalyst (Sb0.67Bi1.33Te3), improving CO2 reduction and oxidation kinetics [4]. A schematic representation of our multiscale approach is depicted in Figure 1. This multiscale approach enables us to make predictions about porosity, deposited species quantities, and CO2 concentration variations-parameters that are difficult to assess through experimental methods alone. Furthermore, it allows us to evaluate these properties as a function of time, providing dynamic insights into the battery’s behavior. This computational methodology can be used to test and screen various battery materials, potentially saving significant time and resources by reducing the need for extensive experimental tests.
@@ 2.2 Ab initio molecular dynamics (AIMD) and Kubo-Greenwood method for electrical conductivity calculation
where E is broadening of the -function. This technical parameter plays an essential role in controlling the algorithm in the Kubo-Greenwood formula, particularly at low frequencies.
@@ 2.4.1 Electrode Reactions
The porous carbon electrode provides a site for the electrochemical reduction of carbon dioxide. During operation, CO2 from the external feed side dissolves in the electrolyte, moves through the pores of the positive electrode, and reacts with lithium ions at the active sites. The reaction considered at the positive electrode is [11]:
where and represent the equilibrium potentials for discharge and charge reactions, respectively [17, 18]. The reaction products (Li2CO3 and C) have limited solubility in the DMSO-based liquid electrolyte (0 mol/L for C and 0.0371 mol/L for Li2CO3). When their concentrations exceed these solubility limits, the electrolyte solution becomes supersaturated [78], leading to the deposition of Li2CO3 and C. This is followed by nucleation and growth of these products in the liquid phase within the porous electrode. These deposited products significantly impact the battery’s performance by covering the active surface area, obstructing pathways for reactive species (Li+ and CO2), and inhibiting further reactions inside the cathode, ultimately contributing to the termination of cell discharge.
@@ 2.4.3 Equations of battery electrochemical properties
denotes the starting porosity of the cathode. It is important to mention that the effect of deposits (Li2CO3 and C) with low conductivity on was not taken into account, and it was assumed that would not change even as the battery reactions occurred.
@@
For modeling the cathode’s electrochemical reaction, a modified version of the Butler-Volmer equation is applied in the model using two rate coefficients. This modification is necessary because the cathode reaction presented in equation 4 depends on both the concentrations of Li+ and CO2, as well as the concentration of Li2CO3/C [84]:
where is local transfer current density between electrode and electrolyte interface, is the molar concentration of species at the cathode surface, and are the anodic and cathodic rate constant, respectively, is a symmetry factor equal to 0.5, is surface or activated overpotential for reaction at cathode, and are the voltage drop and the electrical resistivity across /C film formation, respectively, is the volume fraction of solid /C, and is the theoretical open-circuit potential for reaction.
@@ 3.1 Electrical conductivity from DFT, AIMD and Kubo-Greenwood formula
Calculated dynamic electrical conductivities for both Sb0.67Bi1.33Te3 and Li2CO3 are shown in Figure 2 (a1) and (b1) for a large frequency range and in Figures 2 (a2) and (b2) for a small frequency range, for different values of E. The static electrical conductivities at zero frequency are calculated by interpolating small frequency data shown in Figures 2 (a2) and (b2). The static electrical conductivities calculated for different values of E show the conductive behavior of Sb0.67Bi1.33Te3 (a conductivity of the order of S/m) and insulative behavior of Li2CO3 (a conductivity of the order of S/m). Our computational analysis reveals a static electrical conductivity of S/m for Sb0.67Bi1.33Te3, indicating its strong electrical conducting properties. This finding aligns with the well-established characteristics of bismuth telluride (Bi2Te3) and its alloys, which are extensively employed in thermoelectric devices due to their good electrical conductivity [85, 86]. The high conductivity can be attributed to the semiconductor nature of Bi2Te3 and the antimony (Sb) doping effect, which augments the carrier concentration [87]. Experimental investigations on Bi2Te3 have reported electrical conductivity values as high as S/m at 323 K [87], further corroborating the computational results. Our calculations reveal that Li2CO3 exhibits a static electrical conductivity of S/m at 300 K, confirming its highly insulating nature. This result closely aligns with experimental measurements of S/m at 393 K [88]. The slight difference between calculated and experimental values is consistent with the expected increase in conductivity at higher temperatures, underscoring the reliability of our computational approach. The insulating nature of Li2CO3 can be attributed to its ionic nature and the strong ionic bonds that hinder electron mobility [89]. In Li2CO3, Li+ ions act as the primary charge carriers rather than electrons, a characteristic typical of ionic compounds. The low conductivity of Li2CO3 presents challenges to battery performance, including high internal resistance and passivation of the electrode surface [90, 91]. To address these challenges, potential solutions include promoting Li2CO3 decomposition through conductive additives or catalysts, designing porous electrodes for increased surface area, or directing the discharge reaction towards products that are easier to decompose (such as Li2C2O4).
@@
To comprehensively understand the formation of Li2CO3 and C on the Sb0.67Bi1.33Te3 catalyst surface, we calculated the Gibbs free energy () changes for various elementary reaction steps on this surface as shown below (steps 1-5) similar to the calculations done in Ref. [4]. Here, we used a slightly bigger supercell (as shown in Figure 2 (a1)) compared to the work done in Ref. [4]. The reaction pathway revealed a sequence of intermediate stages, beginning with the adsorption of CO on a Bi site with a Bi-O distance of 3.5 Å, which was marginally endergonic ( eV). The coupled addition of Li+ and an electron to the adsorbed CO formed the LiCO intermediate, a highly exergonic step ( eV). Subsequent interactions between two LiCO intermediates produced Li2CO and CO, with a of eV. Further, the conversion of Li2CO and CO into Li2CO3 and CO∗ ( eV) marked a pivotal step in the pathway. Finally, the generation of carbon was facilitated through the reaction of CO and CO2, an energetically uphill step ( eV) that also produced CO on the surface. These results, illustrated in Figure 2 (c), highlight the essential role of coupled electron-cation transfer processes in the growth of discharge products. The results in this paper are consistent with the work reported in Ref. [4]. The Gibbs free energy of the reaction was calculated using the following equation:
@@
To level up the electrochemical step, a potential was applied to the formation of LiCO2. As can be seen in Figure 2 (c), this potential occurs at 2.3 V versus Li/Li+ electrode [4].
@@
The strong exergonic nature of steps 2-4 suggests that these processes occur readily, while the endergonic nature of the final carbon formation step aligns with the observed preferential adsorption of carbon on the Sb0.67Bi1.33Te3 catalyst surface. The Li2CO3 is formed in step 4, but this step also releases CO∗ and a free surface site (∗). This suggests that Li2CO3 formation occurs near, but not directly on, the catalyst surface, reducing the cathode porosity and possibly clogging the cathode pores.
@@
This section presents the results of MD simulations that elucidate the atomic-level dynamic and solvation properties of the Li-CO2 battery electrolyte depicted in Figure 3 (a). These simulations provide microscopic insights into the electrolyte’s behavior, bridging the gap between the atomic and macroscopic transport properties in our multiscale approach.